# Diff3DS: Generating View-Consistent 3D Sketch via Differentiable Curve Rendering

### v1

@@ 5.1 Text-to-3D Sketch
Pioneering works [28, 41] propose the Score Distillation Sampling (SDS) algorithm, utilizing the prior of a pre-trained 2D text-to-image model to optimize 3D representations. By integrating SDS with Diff3DS, we distill the diffusion prior to generate the 3D sketch $\tilde{\Theta}$ from the input text. Given a pre-trained diffusion model $\epsilon_{\phi}$, text embedding $y$, rendered image $I$ with our rasterizer $\mathcal{R}$ and the noise timestep $t$, image $I$ will be added with noise to obtain a noisy image $I_{t}$: $I_{t}=\sqrt{\bar{\alpha}_{t}}I+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ where $\epsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ is the cumulative product of scaling at timestep $t$. Then the gradient of SDS loss on $\tilde{\Theta}$ is given by:
@@
where $\omega(t)$ is a weight based on the timestep $t$, and the predicted noise sampled by classifier-free guidance (CFG) [13] with weight $\lambda$ denotes as:
@@
Similar with [28], due to the bias of 2D text-to-image model, we encounter serious view inconsistency problem known as the Janus problem. This problem persists even when directional prompts (e.g., "the front view of…") are used to distinguish between different views. To mitigate this issue, we integrate MVDream [35] with our framework, which is able to generate 3D-aware four view images using input text and camera poses.
@@ 5.2 Image-to-3D Sketch
Zero-1-to-3 [22] proposes an image-guided SDS method that conditions on the input view image $\tilde{I}$ and relative camera extrinsic ($R$, $T$), based on its view-conditioned model. Our framework is able to generate 3D sketch $\tilde{\Theta}$ from the input reference image $\tilde{I}$ by distilling the 3D consistent prior:
@@ 5.4 Time Annealing Schedule
To further enhance the performance of the results, we employ a time annealing schedule similar with [43, 57, 35, 14]. During the optimization process, we gradually adjust the maximum and minimum time step for SDS in a linear manner. The large noise timestep focuses on aligning the curves with the semantic content of the input text or image during the initial training phase. Conversely, the small noise timestep employed during the later training phase focuses on further enhancing the details.
@@ 6.1 Implementation Details
We implement our rendering framework in C++/CUDA with a PyTorch interface [27], and follow the STEP standard (ISO 10303-21) [23] to format and save our 3D curve results. In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. Notice that Stable-Zero123 abnegates the relative distance for simplifying, we fix its radius at $2.0$. For the pre-trained model, we apply Stable Diffusion 2.1 [2] and MVDream [35] for the text-to-3D sketch task, with the CFG weight of $50$. And we apply Zero-1-to-3 [22] and Stable-Zero123 [3] for the image-to-3D sketch task, with the CFG weight of $7.5$ followed [22, 3] . For all tasks, the total number of training steps is 4000. Starting from step 2000, we dynamically delete the noise every 100 steps. The training process requires 1 hour for the text-to-3D sketch task and 2 hours for the image-to-3D sketch task on a single NVIDIA A10 GPU with a batch size of 4.
@@ 6.2 Text-to-3D Sketch
To the best of our knowledge, we are the first text-to-3D sketch method. We select the existing text-to-3D object methods DreamGaussian [37] and MVDream [35] as compared baselines, and further choose the text-to-2D sketch method DiffSketcher [49] as the additional reference.
@@
We collect 35 text prompts from previous works [28, 35] and websites. To encourage baseline methods to generate sketch-style results, we append the suffix "sketch in black and white, line drawing" to each prompt inputted to the baselines. We measure the CLIP [32] text-image similarity (CLIP-ScoreT) metric to evaluate the semantic coherence between the input text prompt and the rendered images. For each generated 3D content, we compute the average metric across four orthogonal view images relative to the input text prompt. Table 2 shows the quantitative results of the text-to-3D sketch task. DiffSketcher achieves the best performance, while our method outperforms all the text-to-3D baselines.
Qualitative Comparisons
Fig. 5 shows the qualitative results of the text-to-3D sketch task. We compared the 3D results of the text-to-3D baseline method with our method, and further provide the 2D results of DiffSketcher. Compared text-to-3D baselines utilize Stable Diffusion [2] as the prior provider but do not fully inherit its capability. They struggle to generate the sketch-style content, even with the additional suffix. Our method and DiffSketcher can generate the 3D sketch and 2D sketch similar with the input text.
Table 1: Quantitative results of the text-to-3D sketch task. The CLIP-ScoreT metric is shown.
Method CLIP-ScoreT $\uparrow$
@@
DiffSketcher [49] 33.64 33.98
DreamGaussian [37] 26.25 25.12
MVDream [35] 28.39 27.08
Our 30.35 30.00
Table 2: Quantitative results of the image-to-3D sketch task. The CLIP-ScoreI metric is shown.
Method CLIP-ScoreI $\uparrow$
@@
CLIPasso [40] 83.79 82.72
NEF [51] 66.69 66.53
Our 67.81 68.54
@@ 6.3 Image-to-3D Sketch
To the best of our knowledge, we are the first to generate 3D sketch from a single image. The closest works to ours are Neural Edge Fields (NEF) [51] and 3Doodle [8], which generate 3D curves from multiview images. However, since 3Doodle has not released the code, we select NEF for comparison. NEF reconstructs 3D curves from the 2D multiview edge maps estimated from PidiNet [36]. We also choose the image-to-2D sketch method CLIPasso [40] as the additional reference.
@@
We collect 25 images generated with Imagine [1], and measure the CLIP-Score [32] visual similarity (CLIP-ScoreI) metric to evaluate the semantic coherence between the input reference image and the rendered images. For each generated 3D content, we render four orthogonal view images to calculate the average metrics with the reference image. Table 2 illustrates that CLIPasso performs the best, while our method outperforms NEF.
Figure 6: Qualitative results of the image-to-3D sketch task. NEF fails to exact reasonable 3D curves from the input image.
Qualitative Comparisons
Fig. 6 shows the qualitative results of the image-to-3D sketch task. As a reconstruction method, NEF fails to extract reasonable 3D curves from a single input image because it lacks the ability to predict the unobserved regions. Our method and CLIPasso can generate the 3D sketch and 2D sketch, respectively, conditioned on the input image.
@@ 6.4 Ablations and Analysis
Figure 7: Illustration of the optimization process.
Optimization Process Visualization
Fig. 7 illustrates the visualization of the training optimization process. The optimization starts with randomly initialized curves. As time progresses, curves gradually optimize to resemble the input text or image. Dynamic Noise Deletion intervenes during the middle stage of training to remove the noise generated during the training process.
@@
Ablation Study of Designed Components
@@ 7 Conclusions
In this paper, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. By employing the recent Score Distillation Sampling (SDS) to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the flexible text or single image input. Extensive experiments have demonstrated the potential of our framework.
Limitations and Future Work
First, Diff3DS inherits the sparse gradient issue from DiffVG and can only optimize continuous parameters. Moreover, our downstream task assumption does not distinguish between the view-independent curve (e.g., feature lines) and the view-dependent curve (e.g., contours of smooth surface boundaries), limiting the expressive capacity of the overall shape of a 3D object, which is extensively discussed in 3Doodle. Incorporating the diverse curve representations of 3Doodle is a promising solution. Also, one future direction is to extend our method for scene-level generation task, such as text-to-3D scene sketch.
@@
    In section D, we provide more high quality sketch results.
@@ Appendix A Implementation Details
In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. The default curve number is set to 56. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. Notice that Stable-Zero123 abnegates the relative distance for simplifying  [^2] https://github.com/threestudio-project/threestudio/issues/360, we fix its radius at $2.0$.
For the pre-trained model, we apply Stable Diffusion 2.1 [2] and MVDream [35] for the text-to-3D sketch task, with the CFG weight of $50$. And we apply Zero-1-to-3 [22] and Stable-Zero123 [3] for the image-to-3D sketch task, with the CFG weight of $7.5$ followed by [22, 3] .
@@
Due to rational Bézier curves are projective invariant, the 2D curve $p(t)$ still is a rational Bézier curve defined by projected control points. Given projected 2D control points
@@ Appendix C Closest Distance between a Point and a Rational Bézier Curve
To solve all real roots of these polynomials, inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method [45]. The Newton-Raphson solver obtains its initial guess for intervals from isolator polynomials [34]. For the polynomial $\rho(t)$ with two adjacent real roots $t_{1}$ and $t_{2}$, given any two lower order other polynomials $b(t)$ and $c(t)$, define
@@
where $\rho^{\prime}(t)$ is the derivative, [34] proofs that $a(t)$ or $b(t)$ must has at least one real root in the closed interval $[t_{0},t_{1}]$. Since $\rho(t_{0})=\rho(t_{1})=0$, we know $a(t_{0})a(t_{1})=b(t_{0})b(t_{1})\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})$. Since $t_{0}$ and $t_{1}$ are roots of $\rho$ and $\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})\leq 0$. Thus either $a(t_{0})a(t_{1})\leq 0$ or $b(t_{0})b(t_{1})\leq 0$. Please see the original paper for discussions on multiple roots.
After solving all the real roots of $a$ and $b$ within the $(0,1)$, the intervals for finding the real roots of $\rho$ are then determined. We can then use bisection and the Newton-Raphson method to find all the roots of $\rho$ within these intervals. The key operation is the selection of the polynomials $b(t)$ and $c(t)$, [34] prefer to find the $b(t)$ and $c(t)$ which follow $\mathrm{Degree}(a)+\mathrm{Degree}(b)=\mathrm{Degree}(\rho)-1$. In the upcoming paragraphs, we will discuss the cases of the 4th order polynomial and the 7th order polynomial in section C.3 and section C.4.

### v2

@@ 4.4 Gradients Back-propagation
Figure 3: We generate the 3D sketch $\tilde{\Theta}$, represented as a set of 3D strokes, from the text or image input. We render the raster image $I$ from the random camera via our differentiable renderer (Sec. 4). Then, the pre-trained diffusion model, conditioned on the input, diffuses the rendering $I$ and predicts the pseudo ground truth␣̂$\hat{I}_{0}$. The discrepancies between $\hat{I}_{0}$ and $I$ are used to update the 3D sketch.
@@ 5.1 Text-to-3D Sketch
Pioneering works (Poole et al., 2023; Wang et al., 2023a) propose the Score Distillation Sampling (SDS) algorithm, utilizing the prior of a pre-trained 2D text-to-image model to optimize 3D representations. By integrating SDS with Diff3DS, we distill the diffusion prior to generate the 3D sketch $\tilde{\Theta}$ from the input text. Given a pre-trained diffusion model $\epsilon_{\phi}$, text embedding $y$, rendered image $I$ with our rasterizer $\mathcal{R}$ and the noise timestep $t$, image $I$ will be added with noise to obtain a noisy image $I_{t}$: $I_{t}=\sqrt{\bar{\alpha}_{t}}I+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ where $\epsilon\sim\mathcal{N}(0,I)$ and $\bar{\alpha}_{t}$ is the cumulative product of scaling at timestep $t$. Then the gradient of SDS loss on $\tilde{\Theta}$ is given by:
@@
where $\omega(t)$ is a weight based on the timestep $t$, and the predicted noise sampled by classifier-free guidance (CFG) (Ho & Salimans, 2021) with weight $\lambda$ denotes as:
@@
Similar with (Poole et al., 2023), due to the bias of 2D text-to-image model, we encounter serious view inconsistency problem known as the Janus problem. This problem persists even when directional prompts (e.g., "the front view of…") are used to distinguish between different views. To mitigate this issue, we integrate MVDream (Shi et al., 2024) with our framework, which is able to generate 3D-aware four view images using input text and camera poses.
@@ 5.2 Image-to-3D Sketch
Zero-1-to-3 (Liu et al., 2023) proposes an image-guided SDS method that conditions on the input view image $\tilde{I}$ and relative camera extrinsic ($R$, $T$), based on its view-conditioned model. Our framework is able to generate 3D sketch $\tilde{\Theta}$ from the input reference image $\tilde{I}$ by distilling the 3D consistent prior:
@@ 5.4 Time Annealing Schedule
To further enhance the performance of the results, we employ a time annealing schedule similar with (Wang et al., 2023b; Zhu et al., 2024; Shi et al., 2024; Huang et al., 2024). During the optimization process, we gradually adjust the maximum and minimum time step for SDS in a linear manner. The large noise timestep focuses on aligning the curves with the semantic content of the input text or image during the initial training phase. Conversely, the small noise timestep employed during the later training phase focuses on further enhancing the details.
@@ 6.1 Implementation Details
We implement our rendering framework in C++/CUDA with a PyTorch interface (Paszke et al., 2019). In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. For the pre-trained model, we apply Stable Diffusion 2.1 (sta, ) and MVDream (Shi et al., 2024) for the text-to-3D sketch task. And we apply Zero-1-to-3 (Liu et al., 2023) and Stable-Zero123 (Sta, 2023) for the image-to-3D sketch task. The training process requires 1 hour for the text-to-3D sketch task and 2 hours for the image-to-3D sketch task on a single NVIDIA A10 GPU. More details can be found in the Appendix A.
@@ 6.2 Text-to-3D Sketch
To the best of our knowledge, we are the first text-to-3D sketch method. We select the existing text-to-3D object methods DreamGaussian (Tang et al., 2024) and MVDream (Shi et al., 2024) as compared baselines, and further choose the text-to-2D sketch method DiffSketcher (Xing et al., 2023) as the additional perceptual reference.
Qualitative Comparisons
Fig. 5 shows the qualitative results of the text-to-3D sketch task. We compared the 3D results of the text-to-3D baseline method with our method, and further provide the 2D results of DiffSketcher. To encourage baseline methods to generate sketch-style results, we append the suffix "sketch in black and white, line drawing" to each prompt inputted to the baselines. From the results, it can be seen, even with the additional suffix, the general text-to-3D baselines DreamGaussian and MVDream cannot generate the sketch-style content. For DiffSketcher, the results are more like sketch draws and only in 2D. In contrast, our method can generate 3D view-consistent sketches with clearly defined curves.
@@
We collect 35 text prompts from previous works (Poole et al., 2023; Shi et al., 2024) and websites. Notably, the accurate evaluation of 3D sketch generation is yet to be resolved due to the absence of ground truth sketches. In this work, we measure the CLIP text-image similarity (Radford et al., 2021) (CLIP-ScoreT) and BLIP-Score (Li et al., 2022) metrics to evaluate the consistency of the rendered views with the input text prompt, following previous work (Xing et al., 2024). For all metrics, we render the 3D sketch into 8 views and compute the metric between each view and the input text prompt, and use the averaged value as the final result. Table 2 shows the evaluation results and our method outperforms all the text-to-3D baselines.
Table 1: Evaluation of text-to-3D sketch task.
Method CLIP-ScoreT $\uparrow$ BLIP-Score $\uparrow$
@@
DreamGaussian 0.2561 0.2453 0.4990
MVDream 0.2763 0.2653 0.4949
Diff3DS (Ours) 0.3046 0.3034 0.5038
Table 2: Evaluation of image-to-3D sketch task.
Method Novel Views Reference View
CLIP-ScoreI $\uparrow$ LPIPS $\downarrow$
@@
NEF 0.6612 0.6605 0.3997
3Doodle 0.6734 0.6746 0.2567
Diff3DS (Ours) 0.6768 0.6846 0.2647
Figure 6: Qualitative results of the image-to-3D sketch task. 3Doodle and NEF fail to extract reasonable 3D curves from a single input image.
@@ 6.3 Image-to-3D Sketch
To the best of our knowledge, we are the first to generate 3D sketch from a single image. The closest works to ours are Neural Edge Fields (NEF) (Ye et al., 2023) and 3Doodle (Choi et al., 2024), which generate 3D curves from multiview images. We also choose the image-to-2D sketch method CLIPasso (Vinker et al., 2022) as the additional perceptual reference.
Qualitative Comparisons
Fig. 6 shows the qualitative results of the image-to-3D sketch task. Note that CLIPasso can only generate 2D sketch and is used as a reference for the results of 2D generation methods. As the reconstruction methods, 3Doodle and NEF attain reasonable results for the reference view, but they fail for other views because they lack the ability to predict the unobserved regions from a single input image. In contrast, our method can generate view-consistent 3D sketches.
@@
We collect 25 images generated with Imagine (Ima, 2023). Follow previous works (Qian et al., 2024; Choi et al., 2024), we use CLIP visual similarity (Radford et al., 2021) (CLIP-ScoreI) metric to measure the abstract semantic similarity across the reference image and the rendered novel views, while employing LPIPS (Zhang et al., 2018) metric in the reference view to measure structural semantic similarity. For the CLIP-ScoreI, we calculate the metric between each of the 8 rendered images and the reference image, and use the average value as the final score. Table 2 reports the evaluation results. On the LPIPS metric, our method significantly outperforms NEF and achieves a score comparable to 3Doodle, which has optimized for the reference view with more losses, including the LPIPS and CLIP loss. On the CLIP-Score metric, our method surpasses all baseline approaches. These results demonstrate the superiority of our method.
User Study
We further conduct a user study to evaluate the overall quality of our image-to-3D sketch results. Specifically, we prepared 25 tasks, each of which is composed of three randomly-ordered 3D sketches generated using three methods: NEF, 3Doodle and Diff3DS.
Figure 7: Bar plots of our user study results. The score is of scale 1-5, the higher the better.
We used the Likert scores as the evaluation metric, with a range from 1 to 5 where a higher score indicates that the generated 3D sketch is more preferred by the users. In each task, the participants were asked to score each 3D sketch result based on the following two questions: (1) How well does the 3D sketch fit the input image, e.g., in terms of keeping the similar shape and structure of the input? (2) How is the quality of the 3D sketch, e.g., whether the 4 rendered views are 3D consistent or whether the sketch is visually pleasing? We distributed questionnaires to 40 participants who are CS, EE and Math students and researchers, and got 1000 valid scores in total. The results are summarized in Fig. 7, which highlights the superiority of our method. More details can be found in the Appendix D.
@@ 6.4 Ablations and Analysis
Analysis of Designed Rasterizer
Figure 9: Analysis of designed rasterizer.
Our rasterizer accurately renders the blended colors of overlapping curves and faithfully reproduces occlusion relationships between curves, and enables seamless integration with other 3D sketch generation methods like 3Doodle. To assess its effectiveness, we conduct an experiment to generate colored 3D sketches from multi-view images using both the original 3Doodle and a variant integrated with our rasterizer. As shown in Fig. 9, the original 3Doodle suffers from noticeable color conflict errors across multiple viewpoints (e.g., the support base, which is farther from the camera, is incorrectly rendered in the foreground, occluding the LEGO excavator). These noticeable color conflicts disrupt the object’s visual coherence and lead to semantic ambiguity. In contrast, our rasterizer consistently reproduces accurate occlusion relationships between objects, providing clearer visual and spatial semantics. Additional results and details can be found in the Sec. D.8.
@@ 7 Conclusions
In this paper, we propose Diff3DS, a novel differentiable rendering framework for generating view-consistent 3D sketch by optimizing 3D parametric curves under various supervisions. By employing the recent Score Distillation Sampling (SDS) to distill prior knowledge from pre-trained 2D image generation model, we achieve 3D sketch generation from the flexible text or single image input. Our proposed rasterizer accurately renders the blended colors in overlapping regions and faithfully reproduces occlusion relationships between curves, showcasing its strong potential for generating colored 3D sketches. One limitation of Diff3DS is it inherits the sparse gradient issue from DiffVG and can only optimize continuous parameters. Also, the initial curve number is set manually to achieve a balance between the approximation accuracy and complexity of the results. Moreover, our current implementation does not distinguish between the view-independent curve (e.g., feature lines) and the view-dependent curve (e.g., contours of smooth surface boundaries), limiting the expressive capacity of the overall shape of a 3D object, which is extensively discussed in 3Doodle. Incorporating the diverse curve representations of 3Doodle to Diff3DS is a promising future work. Also, one future direction is to extend our framework for scene-level generation task, such as text-to-3D scene sketch.
Acknowledgement
This work was supported in part by National Natural Science Foundation of China (62202199).
@@ Acknowledgement
    In section D, we provide more analysis and experiment results.
  • $\bullet$
    In section E, we provide more high quality sketch results.
@@ Appendix A Implementation Details
We implement our rendering framework in C++/CUDA with the PyTorch interface, and follow the STEP standard (ISO 10303-21) (Loffredo, 1999) to format and save our 3D curve results. In the experiment, a user-specified number of curves will be randomly initialized within a sphere of radius 1.5. The default curve number is set to 56. We randomly sample the camera position using the radius from 1.8 to 2.0, with the azimuth in the range of -180 to 180 degrees, the elevation in the range of 0 to 30 degrees and the field of view (fov) of 60 degrees. Notice that Stable-Zero123 abnegates the relative distance for simplifying  [^2] https://github.com/threestudio-project/threestudio/issues/360, we fix its radius at $2.0$.
For the pre-trained model, we apply Stable Diffusion 2.1 (sta, ) and MVDream (Shi et al., 2024) for the text-to-3D sketch task, with the CFG weight of $50$. And we apply Zero-1-to-3 (Liu et al., 2023) and Stable-Zero123 (Sta, 2023) for the image-to-3D sketch task, with the CFG weight of $7.5$ followed by (Liu et al., 2023; Sta, 2023) .
@@
For the Dynamic Noise Deletion, only the curves with lengths below a specified threshold will be considered as noise and removed dynamically. The trade-off between noise removal and detail preservation can be balanced by adjusting the threshold. Currently, the threshold is empirically set to 0.1 which is below the $10\%$ of the average curve length.
@@
Due to rational Bézier curves are projective invariant, the projected 2D curve $p(t)$ still is identical to the rational Bézier curve defined by projected control points. Given projected 2D control points
@@ Appendix C Closest Distance between a Point and a Rational Bézier Curve
To solve all real roots of these polynomials, inspired by DiffVG, we solve the polynomial using bisection and the Newton-Raphson method (Press et al., 2007). The Newton-Raphson solver obtains its initial guess for intervals from isolator polynomials (Sederberg & Chang., 1994). For the polynomial $\rho(t)$ with two adjacent real roots $t_{1}$ and $t_{2}$, given any two lower order other polynomials $b(t)$ and $c(t)$, define
@@
where $\rho^{\prime}(t)$ is the derivative, (Sederberg & Chang., 1994) proofs that $a(t)$ or $b(t)$ must has at least one real root in the closed interval $[t_{0},t_{1}]$. Since $\rho(t_{0})=\rho(t_{1})=0$, we know $a(t_{0})a(t_{1})=b(t_{0})b(t_{1})\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})$. Since $t_{0}$ and $t_{1}$ are roots of $\rho$ and $\rho^{\prime}(t_{0})\rho^{\prime}(t_{1})\leq 0$. Thus either $a(t_{0})a(t_{1})\leq 0$ or $b(t_{0})b(t_{1})\leq 0$. Please see the original paper for discussions on multiple roots.
After solving all the real roots of $a$ and $b$ within the $(0,1)$, the intervals for finding the real roots of $\rho$ are then determined. We can then use bisection and the Newton-Raphson method to find all the roots of $\rho$ within these intervals. The key operation is the selection of the polynomials $b(t)$ and $c(t)$, (Sederberg & Chang., 1994) prefer to find the $b(t)$ and $c(t)$ which follow $\mathrm{Degree}(a)+\mathrm{Degree}(b)=\mathrm{Degree}(\rho)-1$. In the upcoming paragraphs, we will discuss the cases of the 4th order polynomial and the 7th order polynomial in section C.3 and section C.4.